OpenStudy (dls):

If f(x)=x^2+lambda x +u be an integral function of the integral variable x then 1)lambda is an integer and u is a rational fraction 2)lambda and u are integers 3)u is an integer and lambda is a rational funcion 4)lambda and u are rational functions

OpenStudy (dls):

@yrelhan4 @shubhamsrg @electrokid

OpenStudy (dls):

\[\Large f(x)=x^2+\lambda x+u\]

OpenStudy (dls):

Since X is an integer variable,lets fix it as 2 now I don't think lambda has to be necessarily integer, lets take lambda as 2.5 so it becomes \[2^2+2 \times 2.5 +u\] still the condition is satisfied,the overall function is yielding integral value. while u has to be integer,I'm not quite sure about my views though ^_^

OpenStudy (dls):

It can't be 1 It can't be 3 It can't be 4 It has to be (2) but If I had an option like lambda CAN be a decimal but u will always be integer I would prefer it or not? However it only implies for rare values,it won't apply for 1.1,1.2..etc..

OpenStudy (yrelhan4):

@shubhamsrg THATS THE BEST I CAN DO. \m/

OpenStudy (dls):

:|

OpenStudy (anonymous):

how about approaching it this way assume "x" and f(x) are rationals and cane be differentiated the \[f(x_0+\Delta x)=f(x_0)+\Delta xf'(x_0)+\frac{\Delta x^2}{2}f''(x_0)\\ f(x_0+\Delta x)=f(x_0)+(2x_0+\lambda)\Delta x+(\Delta x)^2\\\;\\ n+1=n+(2+\lambda)+1\implies \lambda=-2\\\;\\ n-1=n+(2+\lambda)+1\implies\lambda=-4 \] hence, we may conclude that both lambda and u must be integers

OpenStudy (anonymous):

but if you have had \[f(n)=an^2+bn+c\] where n and f(n) are integers, then all of them must be rational functions.